# Artificial Intelligence in Industry 4.0: A Review of Integration Challenges for Industrial Systems

### v1

@@ 1 Literature Search Methodology
Our literature review was conducted via academic search platforms including Google Scholar, Semantic Scholar, and Zeta Alpha. The keywords used were "challenges," "artificial intelligence," "machine learning," "engineering," "Industry 4.0," and "cyber-physical systems." We complemented this search by exploring relevant standards on the websites of the International Organization for Standardization (ISO) and the German Institute for Standardization (DIN). We also reviewed cross-references in key papers, expanding our collection to include critical works that might otherwise have been missed. Our team’s expertise, informed by ongoing AI projects, helped refine our search to include studies that provide deep insights into specific challenges.
@@ 2 Challenges
      Privacy: Many AI models operate in the cloud rather than on local networks, necessitating the transfer of sensitive data to servers potentially located outside Europe. This situation raises concerns about the unauthorized use of data without the company’s consent [25]. Furthermore, legal obligations, such as those imposed by the General Data Protection Regulation (GDPR), further complicate data management and compliance [19, 4].
@@ 3 Discussion
This survey reviewed recent literature to identify the current challenges relevant to contemporary industrial applications. As AI technology evolves, new challenges have emerged that have not been recognized by earlier sources, as illustrated in Figure 3. The proportion of literature addressing and mentioning the identified challenges has increased noticeably over the years, indicating heightened attention. It is important to note that there were relatively few papers in the early years, as shown in Figure 2. Additionally, the frequency with which various challenges are mentioned and addressed in the literature varies considerably, as shown in Figure 4. The most frequently cited challenges primarily fall into two main categories: data and trustworthy AI.
Data is fundamental to the development and performance of AI systems. In the reviewed literature, data quality and data availability are among the most often addressed or mentioned challenges. Many CPS datasets are derived from legacy systems not initially intended for AI data collection [10, 23, 24]. Data latency issues and sensor malfunctions then degrade the data quality [26, 27]. Addressing these challenges often requires extensive data preprocessing or significant hardware upgrades. Various strategies have been proposed to mitigate these issues, including synthetic data generation through simulations [43] and federated learning [44], which allows for the utilization of data across decentralized systems while managing privacy concerns.
Trustworthy AI is increasingly important as AI becomes integrated into more systems. The most addressed or mentioned challenges in this area are robustness, interpretability and quality assurance. These issues are especially critical in Industry 4.0, due to stringent regulatory environments governing technical systems. New regulatory frameworks for AI, such as the AI Act in the European Union [45] and similar proposals in the United States [46] and the United Kingdom [47], underscore the urgency of these concerns. The lack of robustness of AI systems makes a thorough quality assurance crucial, yet the lack of interpretabilityand the dependency on data complicate testing efforts. Approaches such as advanced testing frameworks [48], explainable AI (XAI) [49], and the development of new process models [19, 50] aim to improve AI system robustness, clarify decision-making processes, and meet regulatory demands, thereby enhancing overall system quality and accountability.
@@
The reviewed literature less frequently addressed system integration and workforce challenges. However, practitioners, especially in SMEs, often highlight challenges such as financial constraints, selecting the appropriate AI approach, and addressing the skilled labor shortage [4, 34]. This discrepancy suggests a gap between academic research and practical industry needs. Addressing these barriers through approaches such as automated machine learning (AutoML) [18, 71], transfer learning [5], and enhanced standardization [41] is thus essential for enabling the successful adoption of AI technologies in the industrial sector. In order to design a holistic solution approach, a comprehensive understanding of all potential challenges is necessary.

### v2

@@ Abstract
© 2024 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
1 Introduction
@@ 2 Literature Search Methodology
Our literature review was conducted via academic search platforms including Google Scholar, Semantic Scholar, and Zeta Alpha. The keywords used were "challenges," "artificial intelligence," "machine learning," "engineering," "Industry 4.0," and "cyber-physical systems." We complemented this search by exploring relevant standards on the websites of the International Organization for Standardization (ISO), the International Electrotechnical Commission (IEC) and the German Institute for Standardization (DIN). We also reviewed cross-references in key papers, expanding our collection to include critical works that might otherwise have been missed. Our team’s expertise, informed by ongoing AI projects, helped refine our search to include studies that provide deep insights into specific challenges.
@@ 3 Challenges
      Privacy: Many AI models operate in the cloud rather than on local networks, necessitating the transfer of sensitive data to external servers. This situation raises concerns about the unauthorized use of data without the company’s consent [25]. Furthermore, legal obligations, such as those imposed by the General Data Protection Regulation (GDPR), further complicate data management and compliance [19, 4].
@@ 4 Discussion
This survey reviewed recent literature to identify the current challenges relevant to contemporary industrial applications. As AI technology evolves, new challenges have emerged that have not been recognized by earlier sources, as illustrated in Figure 3. The increasing proportion of literature addressing and mentioning the identified challenges suggests a growing academic interest in these challenges, with earlier years featuring a more limited number of studies, as Figure 2 demonstrates. Notably, the frequency with which various challenges are mentioned and addressed in the literature varies considerably, as shown in Figure 4. The most frequently cited challenges primarily fall into two main categories: data and trustworthy AI.
Data is fundamental to the development and performance of AI systems. In the reviewed literature, data quality and data availability are among the most often addressed or mentioned challenges. Many CPS datasets are derived from legacy systems not initially intended for AI data collection [10, 23, 24]. Data latency issues and sensor malfunctions then degrade the data quality [26, 27]. Addressing these challenges often requires extensive data preprocessing or significant hardware upgrades. Various strategies have been proposed to mitigate these issues, including synthetic data generation through simulations [43, 44] and federated learning [45, 46], which allows for the utilization of data across decentralized systems while managing privacy concerns.
Trustworthy AI is increasingly important as AI becomes integrated into more systems. The most addressed or mentioned challenges in this area are robustness, interpretability and quality assurance. These issues are especially critical in Industry 4.0, due to stringent regulatory environments governing technical systems. New regulatory frameworks for AI, such as the AI Act in the European Union [68] and similar proposals in the United States [69] and the United Kingdom [70], underscore the urgency of these concerns. The lack of robustness of AI systems makes a thorough quality assurance crucial, yet the lack of interpretability and the dependency on data complicate testing efforts. Approaches such as advanced testing frameworks [71, 72], explainable AI (XAI) [73, 74], and the development of new process models [19, 61] aim to improve AI system robustness, clarify decision-making processes, and meet regulatory demands, thereby enhancing overall system quality and accountability.
@@
The reviewed literature less frequently addressed system integration and workforce challenges. However, practitioners, especially in SMEs, often highlight challenges such as financial constraints, selecting the appropriate AI approach, and addressing the skilled labor shortage [4, 34]. This discrepancy suggests a gap between academic research and practical industry needs. Addressing these barriers through approaches such as automated machine learning (AutoML) [18, 75], transfer learning [76, 5], and enhanced standardization [41] is thus essential for enabling the successful adoption of AI technologies in the industrial sector. In order to design a holistic solution approach, a comprehensive understanding of all potential challenges is necessary.